OpenStudy (anonymous):

5t − 7(3 + t) how would i distribute

OpenStudy (anonymous):

xt - w(v + t) = xt - wv - wt = xt - wt - wv = (x - w)t - wv. Just substitute. x=5, w=7, v=3

OpenStudy (anonymous):

ty